# Physical aspects of symmetry breaking in an ideal Bose gas at thermal equilibrium

### v2

@@ III.1 Broken aggregate phase gauge symmetry
The concept of broken gauge symmetry in solids was originally introduced to understand and describe the occurrence of distinct phases between a critical temperature below which solids undergo a certain phase transition as characterized by an order parameter. In the present theory, such phase transitions in atomic gases can be modeled by describing the transition of a mixed wave field $\psi=\psi_{0}\cup\psi_{\perp}$ without coherences and a relative reference frame between condensate and non-condensate components of the Bose gas (that leads to non-zero field distribution in the complex plane) to a coherently interfering condensate and non-condensate quantum field that is gauged to zero average (vanishing) values below the critical temperature for Bose-Einstein condensation. In particular, below the critical temperature, the two sub-components condensate and non-condensate are coherently coupled within a certain coherence time (at each realization of the quantum field).
Different from standard approaches known so far, within the present theory, it is, in particular, possible to explain that it is not the entire gauge field symmetry that is broken, but only the relative symmetry between condensate and non-condensate fields that are coherently interacting below the critical temperature. In contrast, in the limit of large temperatures, the two components do not build a relative reference frame for symmetry breaking but are independently mixed, and therefore the total quantum field builds a ball-shaped distribution in complex space. For finite temperatures, there is theoretically no non-zero coherence, however, different field components only interfere destructively, if the coherence time is larger than the time uncertainty. This defines a physical transition for symmetry breaking when the coherence time exceeds this (minimal) uncertainty threshold.
Please note that this is the case only in three-dimensional setups, since in lower dimensions the single-particle wave functions cannot be assumed to obey a non-zero average. Thus, for reduced dimensionality, the assumption that $\psi=0$ doesn’t remain valid, i.e. one has to assume a general function $\psi=\psi(N-N_{0})$. The broken phase gauge symmetry of the (left-shifted) condensate quantum field is illustrated in the lower right section of Fig. 2. As compared to the star-shaped quantum field in Fig. 3, it is visible that the (relative) symmetry is broken concerning the zero point axis assuming thermal equilibrium.
@@ III.2 Broken gauge symmetry from phase correlations
As indicated from the derivation of Eq. (8), the time evolution of the quantum field is one directional in units of the time scale $\tau_{0}$. It is important to recall that the concept of one-directional time evolution can also be derived ab initio from the following numerical calculus. Considering the correlation function in Eq. (14), one may straightforwardly verify that the initial phase of two (initially) uncorrelated quantum fields is distributed (quantized) around multiples of $2\pi$, see Fig. 5. As discussed before in Ref. ref-7 , the correlations around zero phase arise from the separate propagation of the independent fields towards the Boltzmann equilibrium, since at each calculation cycle, the two quantum fields are numerically calculated ab initio and independently. The latter fact ensures that correlations around zero phase can not be only caused by the interference of pairwise independent particle matter waves, since the phases are not purely random, but pre-defined by the affinity of the wave fields to approach the Boltzmann equilibrium that leads to forward time propagation with similar distributions of partial phases (i. e. effectively zero relative phases).
@@
where $\omega m({\bf{k}})$ is the distribution of modes among the energy levels $\bf{k}$, one may conclude that the adjoint quantum fields interfere at multiples of the time scale $\tau_{0}$. Please note that from the symmetry of the total wave field concerning the time scale $\tau_{0}$, i. e. $\psi(t+\tau_{0})=\psi(t)$, the spectrum of the spatially averaged quantum field (order parameter) of the theory remains preserved, as shown in Fig. 4 for the same model parameters as in Fig. 2. Numerically, the manifold of the total wave field at the Boltzmann equilibrium with $\mu\rightarrow 0^{+}$ can simply be extracted from the boundary condition that $\phi=\pm\epsilon\pi$, where the parameter $\epsilon$ ideally tends to zero and has been chosen $0.05$ in the numerical simulation shown in Fig. 6.
@@ III.3 Direction of time from constructive interference
As we understand from the present model, the direction of time manifests itself as an interference process of forward and backward propagating wave fields $\psi_{+}$ and $\psi_{-}$. Considering only quantum field states with complex number representations within the shown manifold, explains directional time as a propagating quantity that arises from the constructive interference of partial matter waves in the Bose gas. The smallest unit time scales for the forward propagation of time is thus defined by the time scales at which partial matter waves interfere (from zero to a few microseconds), whereas, from destructive interference (dynamical localization in complex space), one may deduce that there exists a certain coherence time of the total field that defines a time scale for the decay of different realizations of the thermal Boltzmann equilibrium in the present parameter regime.
@@
Above the critical temperature, the real part of the fugacity spectrum shows a gap between the ring of constant absolute time, indicating that the symmetry of the quantum field can, in principle, not be broken without externally induced quantum fluctuations of energy and the corresponding entropy. Decreasing temperature by external cooling to decrease the gap between the outer ring and the inner linear (symmetry breaking) part of the fugacity spectrum finally leads to a gapless fugacity spectrum. However, due to the uncertainty of the time variable, there is also a finite coupling probability between symmetric and asymmetric parts of the fugacity spectrum, in the parameter range, where the fugacity spectrum obeys a gap between the two principally unconnected spectra close above the critical temperature, which leads to spontaneous symmetry breaking, see Fig. 8.
The probabilistic modeling of the dilute Bose gas of $N$ particles below the critical temperature is described by the equation of conditional probabilities
@@
in the present numerical framework, which defines the relative probability of a quantum particle to switch from a state $1$ with a corresponding chemical potential $\mu_{1}$ to a next state $2$ with a chemical potential $\mu_{2}$ for a Bose gas at thermal equilibrium, where the chemical potential is defined by the intrinsic particle number conserving equation.
Applying only the constraint of particle number conservation as defined by the conservation equation in Eq. (9), numerical sampling leads to a distribution of typical scales for $\Gamma_{1}t$ (oscillation time) and $\Gamma_{2}t$ (coherence time), shown in Fig. 7. From the numerical sampling, it is observed that $\Gamma_{1}t$ is proportional to the real part of the quantum phase and distributed around zero measure in the given parameter range (as defined in Fig. 7 - upper figure). The real part $\Gamma_{1}t$ defines the average oscillation frequency, which means that the quantum particle (with constant particle number at finite temperature) quickly couples to excited single particle quantum states, related to $\Gamma_{1}t\sim 0$ in the form of a standing wave. The distribution of the phase $\Gamma_{2}t$, that is, the effective coherence time for the particle weighted with the decay constant $\Gamma_{2}$, is shown in the lower figure of Fig. 7. The typical range of the time scale $\Gamma^{-1}_{2}$ is between $300~{}\mu s$ and $450~{}\mu s$ (see Figs. 7 and 8) in the present setup. To estimate the time scale in a physical framework, time is scaled in terms of the quantity $\tau=2\pi\times(\omega_{x}\omega_{y}\omega_{z})^{-1/3}$, since any frequency of the equilibrated system below that value is effectively zero in the thermally stable quantum system, because of the finite energy uncertainty. From the present model, it is possible to, in particular, conclude that the so-derived time variable of a bosonic quantum field can provide a fundamental quantization of time only on the scale of quantized time steps $\tau_{0}$ defined by different realizations of the Boltzmann equilibrium. Time intervals smaller than the value of $\tau_{0}$ randomly vary from effectively zero to tens of microseconds at different realizations of the quantum field, which leads to a quasi-continuous numerical time variable without a fundamental unit time (i.e. varying time step sizes) for time steps smaller than the time scale $\tau_{0}$.
@@ IV Discussion
The fact that the conjugate equation precisely describes the backward time evolution of the adjoint quantum states is called the time-reversal symmetry of the Schrödinger equation. Very weakly interacting Bose gases that can be modeled and described in terms of an ideal Bose gas model build a fundamental theoretical system to study quantum effects such as spontaneous symmetry breaking on a measurable scale. The corresponding eigenstates and eigenenergies are not only straightforwardly defined and derivable formally, but can also be calculated numerically in terms of three-dimensional harmonic oscillator states. From the definition of the quantum field in Eq. (7), which is a spatially integrated representation of the quantum field in the complex number plane, it is possible to illustrate that Glauber coherent states built the specific solutions of the Schrödinger equation for an ideal Bose gas with the complex variables $\alpha=|\psi|{\rm e}^{i\phi}$ that define the corresponding symmetric gauge fields as a function of the phase variable $\phi$.
In a local representation of the quantum field, in very generality and mathematical and formal terms, the local symmetry of the quantum system can formally only be broken from a formal process that either modifies the fundamental equation of state or the corresponding gauge fields, respectively, called gauge fixing ref-16 . In ideal Bose gases, such formal projection can, therefore, not only be understood by picking out a certain phase variable by a spooky unphysical action, such as the definition of a mathematical variable by an external observer. As shown in the sequel of the present theory, following the ansatz of a non-local order parameter, the occurrence of spontaneously broken gauge fields can be ideally explained in physical terms by assuming the convergence of the (independent) quantum fields to the Boltzmann equilibrium with a vanishing total chemical potential at constant particle number and temperature of the Bose gas. Thus, the symmetry of the considered wave field is spontaneously broken not from defining a specific formal (zero) gauge of the wave field, but from the physical inset of quantum coherence at the thermal equilibrium, while the underlying total symmetry of the wave field in complex number representation remains preserved. Finally, the one-directional propagation of time is a direct formal consequence of the symmetry breaking of the Schrödinger equation from the assumption of a Boltzmann equilibrium.
@@ V Conclusion
It is particularly interesting to formulate the present quantum field description in pioneering future works on dynamical localization in other systems, such as neuronal networks out of thermal equilibrium in a completely different parameter regime.

### v4

@@ III.1 Broken aggregate phase gauge symmetry
The concept of broken gauge symmetry in solids was originally introduced to understand and describe the occurrence of distinct phases between a critical temperature below which solids undergo a certain phase transition, as characterized by an order parameter. In the present theory, such phase transitions in atomic gases can be modeled by describing the transition of a mixed wave field $\psi=\psi_{0}\cup\psi_{\perp}$ without coherences and a relative reference frame between condensate and non-condensate components of the Bose gas (that leads to a non-zero field distribution in the complex plane) to a coherently interfering condensate and non-condensate quantum field that is gauged to zero average (vanishing) values below the critical temperature for Bose-Einstein condensation. In particular, below the critical temperature, the two sub-components, condensate and non-condensate, are coherently coupled within a certain coherence time (at each realization of the quantum field).
Different from standard approaches known so far, within the present theory, it is, in particular, possible to explain that it is not the entire gauge field symmetry that is broken, but only the relative symmetry between condensate and non-condensate fields that are coherently interacting below the critical temperature. In contrast, in the limit of large temperatures, the two components do not build a relative reference frame for symmetry breaking but are independently mixed, and therefore, the total quantum field builds a ball-shaped distribution in complex space. For finite temperatures, there is theoretically no non-zero coherence; however, different field components only interfere destructively, if the coherence time is larger than the time uncertainty. This defines a physical transition for symmetry breaking when the coherence time exceeds this (minimal) uncertainty threshold.
Please note that this is the case only in three-dimensional setups, since in lower dimensions the single-particle wave functions cannot be assumed to obey a non-zero average. Thus, for reduced dimensionality, the assumption that $\psi=0$ doesn’t remain valid, i.e., one has to assume a general function $\psi=\psi(N-N_{0})$. The broken phase gauge symmetry of the (left-shifted) condensate quantum field is illustrated in the lower right section of Fig. 2. As compared to the star-shaped quantum field in Fig. 3, it is visible that the (relative) symmetry is broken concerning the zero point axis, assuming thermal equilibrium.
@@ III.2 Broken gauge symmetry from phase correlations
As indicated from the derivation of Eq. (8), the time evolution of the quantum field is one-directional in units of the time scale $\tau_{0}$. It is important to recall that the concept of one-directional time evolution can also be derived ab initio from the following numerical calculus. Considering the correlation function in Eq. (14), one may straightforwardly verify that the initial phase of two (initially) uncorrelated quantum fields is distributed (quantized) around multiples of $2\pi$, see Fig. 5. As discussed before in Ref. ref-7 , the correlations around zero phase arise from the separate propagation of the independent fields towards the Boltzmann equilibrium, since at each calculation cycle, the two quantum fields are numerically calculated ab initio and independently. The latter fact ensures that correlations around zero phase can not be only caused by the interference of pairwise independent particle matter waves, since the phases are not purely random, but pre-defined by the affinity of the wave fields to approach the Boltzmann equilibrium that leads to forward time propagation with similar distributions of partial phases (i. e. effectively zero relative phases).
@@
where $\omega m({\bf{k}})$ is the distribution of modes among the energy levels $\bf{k}$, one may conclude that the adjoint quantum fields interfere at multiples of the time scale $\tau_{0}$. Please note that from the symmetry of the total wave field concerning the time scale $\tau_{0}$, i.e.. $\psi(t+\tau_{0})=\psi(t)$, the spectrum of the spatially averaged quantum field (order parameter) of the theory remains preserved, as shown in Fig. 4 for the same model parameters as in Fig. 2. Numerically, the manifold of the total wave field at the Boltzmann equilibrium with $\mu\rightarrow 0^{+}$ can simply be extracted from the boundary condition that $\phi=\pm\epsilon\pi$, where the parameter $\epsilon$ ideally tends to zero and has been chosen $0.05$ in the numerical simulation shown in Fig. 6.
@@ III.3 Direction of time from constructive interference
As we understand from the present model, the direction of time manifests itself as an interference process of forward and backward propagating wave fields $\psi_{+}$ and $\psi_{-}$. Considering only quantum field states with complex number representations within the shown manifold, directional time is a propagating quantity that arises from the constructive interference of partial matter waves in the Bose gas. The smallest unit time scales for the forward propagation of time are thus defined by the time scales at which partial matter waves interfere (from zero to a few microseconds), whereas, from destructive interference (dynamical localization in complex space), one may deduce that there exists a certain coherence time of the total field that defines a time scale for the decay of different realizations of the thermal Boltzmann equilibrium in the present parameter regime.
@@
Above the critical temperature, the real part of the fugacity spectrum shows a gap between the ring of constant absolute time, indicating that the symmetry of the quantum field can, in principle, not be broken without externally induced quantum fluctuations of energy and the corresponding entropy. Decreasing temperature by external cooling to decrease the gap between the outer ring and the inner linear (symmetry-breaking) part of the fugacity spectrum finally leads to a gapless fugacity spectrum. However, due to the uncertainty of the time variable, there is also a finite coupling probability between symmetric and asymmetric parts of the fugacity spectrum, in the parameter range, where the fugacity spectrum obeys a gap between the two principally unconnected spectra close above the critical temperature, which leads to spontaneous symmetry breaking, see Fig. 8.
The probabilistic modeling of the dilute Bose gas with an average of $N$ particles below the critical temperature is described by the equation of conditional probabilities
@@
in the present numerical framework, which defines the relative probability of a quantum particle to switch from a state $1$ with a corresponding chemical potential $\mu_{1}$ to a next state $2$ with a chemical potential $\mu_{2}$ for a Bose gas at thermal equilibrium, where the chemical potential is defined by the intrinsic particle number conserving equation for the average number of particles.
Applying only the constraint of particle number conservation as defined by the conservation equation in Eq. (9), numerical sampling leads to a distribution of typical scales for $\Gamma_{1}t$ (oscillation time) and $\Gamma_{2}t$ (coherence time), shown in Fig. 7. From the numerical sampling, it is observed that $\Gamma_{1}t$ is proportional to the real part of the quantum phase and distributed around zero measure in the given parameter range (as defined in Fig. 7 - upper figure). The real part $\Gamma_{1}t$ defines the average oscillation frequency, which means that the quantum particle (with constant particle number at finite temperature) quickly couples to excited single particle quantum states, related to $\Gamma_{1}t\sim 0$ in the form of a standing wave. The distribution of the phase $\Gamma_{2}t$, that is, the effective coherence time for the particle weighted with the decay constant $\Gamma_{2}$, is shown in the lower figure of Fig. 7. The typical range of the time scale $\Gamma^{-1}_{2}$ is between $300~{}\mu s$ and $450~{}\mu s$ (see Figs. 7 and 8) in the present setup. To estimate the time scale in a physical framework, time is scaled in terms of the quantity $\tau=2\pi\times(\omega_{x}\omega_{y}\omega_{z})^{-1/3}$, since any frequency of the equilibrated system below that value is effectively zero in the thermally stable quantum system, because of the finite energy uncertainty. From the present model, it is possible to, in particular, conclude that the so-derived time variable of a bosonic quantum field can provide a fundamental quantization of time only on the scale of quantized time steps $\tau_{0}$ defined by different realizations of the Boltzmann equilibrium. Time intervals smaller than the value of $\tau_{0}$ randomly vary from effectively zero to tens of microseconds at different realizations of the quantum field, which leads to a quasi-continuous numerical time variable without a fundamental unit time (i.e., varying time step sizes) for time steps smaller than the time scale $\tau_{0}$.
@@ IV Discussion
The fact that the conjugate equation precisely describes the backward time evolution of the adjoint quantum states is called the time-reversal symmetry of the Schrödinger equation. Very weakly interacting Bose gases that can be modeled and described in terms of a non-interacting Bose gas model build a fundamental theoretical system to study quantum effects such as spontaneous symmetry breaking on a measurable scale. The corresponding eigenstates and eigenenergies are not only straightforwardly defined and derivable formally, but can also be calculated numerically in terms of three-dimensional harmonic oscillator states. From the definition of the quantum field in Eq. (7), which is a spatially integrated representation of the quantum field in the complex number plane, it is possible to illustrate that Glauber coherent states are built the specific solutions of the Schrödinger equation for an ideal Bose gas with the complex variables $\alpha=|\psi|{\rm e}^{i\phi}$ that define the corresponding symmetric gauge fields as a function of the phase variable $\phi$.
In a local representation of the quantum field, in very generality and mathematical and formal terms, the local symmetry of the quantum system can formally only be broken from a formal process that either modifies the fundamental equation of state or the corresponding gauge fields, respectively, called gauge fixing ref-16 . In non-interacting Bose gases, such formal projection can, therefore, not only be understood by picking out a certain phase variable by a spooky unphysical action, such as the definition of a mathematical variable by an external observer. As shown in the sequel of the present theory, following the ansatz of a non-local order parameter, the occurrence of spontaneously broken gauge fields can be ideally explained in physical terms by assuming the convergence of the (independent) quantum fields to the Boltzmann equilibrium with a vanishing total chemical potential at constant particle number and temperature of the Bose gas. Thus, the symmetry of the considered wave field is spontaneously broken not from defining a specific formal (zero) gauge of the wave field, but from the physical inset of quantum coherence at the thermal equilibrium, while the underlying total symmetry of the wave field in complex number representation remains preserved. Finally, the one-directional propagation of time is a direct formal consequence of the symmetry breaking of the Schrödinger equation from the assumption of a Boltzmann equilibrium.
@@ V Conclusion
It is particularly interesting to formulate the present quantum field description in pioneering future works on dynamical localization in other systems, such as neuronal networks, out of thermal equilibrium in a completely different parameter regime.